OpenStudy (anonymous):

Is the LCD of this equation -5/(3x-9)+4/(x-3)=5/6 18(x-3) or 6(x-3)

OpenStudy (anonymous):

This should clear it up a little bit \[\frac{-5}{(3x-9)}+\frac{4}{(x-3)}=\frac{5}{6}\]

OpenStudy (anonymous):

3*x-9

OpenStudy (anonymous):

Am still lost because when 3x-9 is factored it is 3(x-3)... and you have to factor the 6 in 5/6 to 3*2 and when you find the common factor in them shouldn't it be 3*2(x-3)...?

OpenStudy (anonymous):

well for the lhs it is 3x-9

OpenStudy (anonymous):

Umm, how did you get that timo?

OpenStudy (anonymous):

but if you set equation = to 0 -5/(3*x-9)+4/(x-3)-5/6=0 lhs common denom is 6*x-18

OpenStudy (anonymous):

Oh, okay I see where your coming from... Idk which one to go with... but thanks so much!!!

OpenStudy (anonymous):

LOL ask around also cuzz I am not exactly sure :P

OpenStudy (anonymous):

okay... thanks

OpenStudy (anonymous):

it also simplifies to this 7/(3*(x-3)) = 5/6

OpenStudy (anonymous):